Angles A and B are supplementary angles because their sum is 180 degrees. The measure of angle A is ten less than 4 times the measure of angle B. Find the measure of angle B.
45 degrees
38 degrees
142 degrees
36 degrees

Answer by richwmiller(17219)

a+b=180
a=-10+4b
plug -10+4b for a in a+b=180

a=142
b=38